OpenStudy (anonymous):

for the graph of r = 1 + 2cos(2θ) on rectangular cordinate, when θ=0 shouldnt r be 5?

OpenStudy (anonymous):

its graph gives a y intercept of 3

OpenStudy (anonymous):

2cos(2θ) (θ=0) = 2(2.1) = 4 ?

OpenStudy (anonymous):

cos(2θ) for θ=0 is 1 so 2*cos(2θ)=2

OpenStudy (anonymous):

cos(2θ)=cos(2*0)=cos(0)=1 so r= 1 + 2cos (2θ) r=1+2*1 =1+2 =3

OpenStudy (imstuck):

No, you would solve it by replacing 0 for the theta. When you multiply 2 times 0 you get 0. So what you are looking for is the cos(0) in radians on your calculator. It comes to 1. So you have r = 1 + 2(1) which is 3.

OpenStudy (imstuck):

See? We all have the same answers! At least we have that! Although we may have come to those answers in different ways with different explanations, they are all the same, so pick whichever way is easiest to understand and go with it!

OpenStudy (anonymous):

so if θ=pi/3 than r would be 2?

OpenStudy (anonymous):

on no i mean 0

OpenStudy (anonymous):

thanks guys!